# OpticalAging: Real-time Presbyopia Simulation for Inclusive Design via Tunable Lenses

### v1

@@ 4.4. Part II: Discussion
4.5. (Part III) Exploratory Case Study: Evaluating Impact on Design Practice
While the previous tasks assessed the simulation’s effectiveness and subjective impact, this exploratory case study investigated its practical application in a design context. The goal was to observe how using the OpticalAging simulation might influence the process of creating and evaluating accessible products.
@@
Participants: Two participants were recruited, a professional product designer and a graduate design student (male, 27 years old). This pairing aimed to capture diverse perspectives, balancing industry experience with an academic viewpoint to understand the tool’s potential applicability across different career stages and design contexts.
@@
Procedure: Participants first underwent the same calibration procedure as in previous tasks, using their personal laptops in a familiar environment (Figure 10). We employed a within-subjects design with two 45-minute conditions, separated by a 15-minute break. The first condition was always Free Design, where participants designed the presbyopia-friendly medicine bottle and label using their preferred methods, without our presbyopia simulation. In the second Simulation Condition, participants used the OpticalAging eyewear (set to 50s mode) to evaluate and redesign their initial concepts. In both phases, participants could use a provided 3D printer and laser printer to prototype their designs. We provided sample label content (medicine name, instructions). Designers considered bottle dimensions, capacity, ergonomics, label layout, typography, and color contrast. The primary goal given to participants was to design a bottle and label optimized for ease of use, particularly focusing on handling ergonomics and information legibility (dosage, warnings, etc.) for individuals likely experiencing presbyopia. While standard aesthetic considerations were relevant, functional accessibility for the target user group was the main priority. Following the tasks, participants completed a semi-structured interview and questionnaire assessing the simulation’s ease of use, impact on design decisions, and overall usefulness.
@@
4.6. Exploratory Case Study: Results
@@
Figure 12. This figure presents the response distribution from two individuals to a baseline question (B) and seven subsequent questions (Q1-Q7) about their experience using the presbyopia simulation method. Rectangles marked with a P represent the feedback from a professional designer, while circles with a S signify the design student’s responses. The baseline query (B) collected initial knowledge on presbyopia prior to the simulation engagement. Questions Q1-Q7 evaluated various aspects of the simulation experience and its impact on the design process aimed at presbyopia users. Responses were recorded on a 7-point Likert scale ranging from 1 (Strongly Disagree) to 7 (Strongly Agree).
@@
Questionnaire Results: Both the professional designer and design student initially showed limited understanding of presbyopia (B), each scoring 2 out of 7. Figure 12 presents the complete results. The post-simulation feedback revealed high effectiveness across various design criteria. The approach was especially beneficial for evaluating and refining visual accessibility elements (Q4), with the professional designer giving it a perfect score of 7/7. Both participants found the method highly effective for selecting suitable typefaces and font styles (Q1), determining text sizes and hierarchy (Q2), and improving the overall design process for presbyopic users (Q5), with ratings between 5 and 7. The simulation also proved valuable for choosing colors (Q3) and seamlessly integrating into existing workflows (Q6). Additionally, both participants expressed a strong inclination to use this approach in future designs for presbyopic users (Q7), with the design student rating it 6/7 and the professional designer 5/7. These findings suggest that our simulation approach greatly improved the designers’ capability to create accessible and user-friendly designs for individuals with presbyopia.
@@
Qualitative Results and Insights: Emphasizing the professional designer’s expertise, the feedback revealed that the most valued feature is the ability to instantly review the current design plan, which can greatly streamline the design iteration process. In the conventional design approach, needs of target users, such as individuals with presbyopia, were treated as a checklist, followed by a prolonged wait for client feedback on each iteration. Viewing from the actual presbyopic user’s perspective enables agile iteration and improves the designer’s confidence. When it comes to design requirements from clients or design handbooks, font size selection often follows vague guidelines like “use a font size larger than or equal to a specified value.” With simulation eyewear, he could refine the font size and style selection within a small range, balancing readability and density of information. The design student observed that, although larger texts had initially been selected, the simulation effectively underscored the necessity of further font size adjustments to better accommodate presbyopic patients. Both designers expressed that the first-person-view simulation enriched their understanding and perception of how everyday objects might appear to their parents and design audiences. This empathetic understanding led to a refined bottle design featuring a larger size, pinchable bottom, and reduced neck and cap diameter for improved dosage control. The design student also highlighted a difference between monocular and binocular viewing, noting that vision seemed much clearer (“as usual”) when looking through the simulation eyewear with only one eye open. This difference could be leveraged to facilitate a smooth transition from normal vision (single-eye viewing) to a presbyopic simulation (dual-eye viewing) without removing the simulation eyewear during design concept evaluations.
4.7. Exploratory Case Study: Discussion
Findings from this small-scale exploratory case study (N=2) should be interpreted cautiously due to the limited sample size. However, the observations hint at several potential areas where the simulation approach might offer value to inclusive design:
@@ 4.7. Exploratory Case Study: Discussion
    Potential Empathetic Design Shift: Observations suggest the simulation could help designers shift from abstract to embodied empathy, potentially aligning with Suri’s ’designing in a state of empathy’ by offering a direct, albeit simulated, user experience (Suri, 2003).
@@
    Indications of Design Process Enhancement: The immediate feedback observed in this case hints at the potential for a more efficient inclusive product development process, though this requires validation in broader contexts.
@@
    Highlighting Potential Theory-Practice Gaps: The designers’ experiences seemed to underscore possible disconnects between abstract accessibility guidelines and the practical challenges faced during design work involving visual impairments.
  4. (4)
    Potential Pathway for Ethical Design: The simulation might offer a supplementary tool for designers, potentially reducing the need for very early-stage testing with vulnerable user groups, aligning with ethical research practices that prioritize user well-being.
  5. (5)
    Possible Educational Implications: The differing perspectives observed between the student and the professional raise questions about potential benefits for bridging knowledge gaps in design education, suggesting an area for future pedagogical research.
  6. (6)
    Avenues for Technological Exploration: The observation regarding monocular viewing suggests further investigation could explore ways to leverage such perceptual differences, potentially leading to improved transitions between vision conditions or integration with AR for more immersive design systems.
Consistent with the exploratory nature of this N=2 study, the findings primarily suggest the tool can potentially aid designers in evaluating specific visual aspects of their designs from a simulated presbyopic perspective. For instance, experiencing difficulty reading smaller fonts firsthand prompted direct modifications in this case. However, designers must remain vigilant against substituting this simulated check for authentic user feedback, recognizing that simulation only captures specific perceptual aspects and cannot replicate the broader context, lived experience, or compensatory strategies employed by individuals with presbyopia.
5. Limitations
Several limitations should be considered when interpreting these findings.
1. Inherent Nature of Simulation: A primary limitation stems from simulation itself. As highlighted by critiques within HCI and disability studies (Bennett and Rosner, 2019; Tigwell, 2021), simulations risk oversimplifying lived experiences, potentially reinforcing stereotypes or substituting superficial understanding for deep user engagement. While OpticalAging simulates a specific physiological change (near-point shift), it cannot replicate the full psychosocial impact, adaptation, or compensatory strategies associated with presbyopia. Its use must be carefully situated within ethical design practices that prioritize user collaboration.
2. Simulation Fidelity: The current implementation presents limitations in its visual representation. Specifically, it renders a uniform blur, which differs from preliminary feedback from older adults describing experiences like ’ghosting’ or ’double-imaging’ at close distances. Furthermore, the hardware currently cannot simulate astigmatism, which often co-occurs with myopia (Kleves, 2021; Fresina et al., 2012). The limited visual field of the eyewear was also noted to cause some visual discomfort and reduced immersion for participants.
3. Study Scope and Design: Our user studies have scope limitations. Simulation Modes: Parts II and III focused solely on the 50s simulation mode. Exploring other modes (e.g., 40s for onset, 60s+ for advanced presbyopia) could yield further insights into the condition’s progression and impact. Tasks and Duration: The exploratory case study involved a single design task (medicine bottle) and a short duration. This limits the generalizability of findings to other product categories or real-world collaborative design processes, which often involve teamwork, extended timelines, and specialized roles (e.g., graphic designers, regulatory experts). The short duration also prevents assessment of the long-term impact of using the tool on design practices or empathy development; longitudinal studies would be needed for this.
4. Case Study Generalizability: The exploratory case study involved only two participants (one student, one professional). While insightful, findings from this small sample cannot be generalized to the broader design community and highlight the need for larger-scale validation.
@@ 6. Conclusion and Future Works
This paper presented OpticalAging, a real-time, optical see-through approach using tunable lenses to simulate the first-person visual perspective of presbyopia. Our findings, blending quantitative near-point validation with qualitative insights from designers, suggest that experiencing simulated presbyopic blur can significantly enhance understanding and empathy, potentially leading to more intuitive and user-centered designs for older adults. By providing a tool to help bridge the gap between abstract accessibility knowledge and embodied appreciation of perceptual challenges, this work offers a step towards more effective age-inclusive design practices.
@@
Ultimately, this research contributes a novel tool intended to foster more compassionate and accessible design. While technology continues to advance, tools like OpticalAging can support designers today in proactively considering age-related visual changes. However, it is crucial to reiterate that OpticalAging should serve as a supplementary tool, not a substitute for direct engagement and collaboration with individuals experiencing presbyopia. Insights gained from simulation must always be validated and contextualized through user-centered research and co-design, ensuring that technology development remains grounded in lived experience.

### v2

@@ 4.5. Part III: An Exploratory Case Study Evaluating Impact on Design Practice
Participants: Two participants (male, 27 years old) were recruited, a professional product designer and a graduate design student. This pairing aimed to capture diverse perspectives, balancing industry experience with an academic viewpoint to understand the tool’s potential applicability across different career stages and design contexts.
@@
Procedure: Participants first underwent the same calibration procedure as in previous tasks, using their personal laptops in a familiar environment (Figure 9). We employed a within-participants design with two 45-minute conditions, separated by a 15-minute break. The first condition was always Free Design, where participants designed the presbyopia-friendly medicine bottle and label using their preferred methods, without our presbyopia simulation. In the second Simulation Condition, participants used the OpticalAging eyewear (set to 50s mode) to evaluate and redesign their initial concepts. In both phases, participants could use a provided 3D printer and laser printer to prototype their designs. We provided sample label content (medicine name, instructions). Designers considered bottle dimensions, capacity, ergonomics, label layout, typography, and color contrast. The primary goal given to participants was to design a bottle and label optimized for ease of use, particularly focusing on handling ergonomics and information legibility (dosage, warnings, etc.) for individuals likely experiencing presbyopia. While standard aesthetic considerations were relevant, functional accessibility for the target user group was the main priority. Following the tasks, participants completed a semi-structured interview and questionnaire assessing the simulation’s ease of use, impact on design decisions, and overall usefulness.
@@
4.6. Part III: Results
@@
Questionnaire Results: Both the professional designer and design student initially showed limited understanding of presbyopia (Baseline), each scoring 2 out of 7. Post-simulation feedback revealed high effectiveness across various design criteria. The approach was especially beneficial for evaluating and refining visual accessibility elements (Q4), with the professional designer giving it a perfect score of 7/7. Both participants found the method highly effective for selecting suitable typefaces (Q1), determining text sizes (Q2), and improving the overall design process (Q5), with all ratings ranging between 5 and 7. The simulation also proved valuable for choosing colors (Q3) and seamlessly integrating into existing workflows (Q6). Additionally, both participants expressed a strong inclination to use this approach in future designs (Q7), with the design student rating it 6/7 and the professional designer 5/7. These findings suggest that our simulation approach greatly improved the designers’ capability to create accessible and user-friendly designs.
@@
Qualitative Results and Insights: Emphasizing the professional designer’s expertise, the feedback revealed that the most valued feature is the ability to instantly review the current design plan, which can greatly streamline the design iteration process. In the conventional design approach, needs of target users, such as individuals with presbyopia, were treated as a checklist, followed by a prolonged wait for client feedback on each iteration. Viewing from the actual presbyopic user’s perspective enables agile iteration and improves the designer’s confidence. When it comes to design requirements from clients or design handbooks, font size selection often follows vague guidelines like “use a font size larger than or equal to a specified value.” With simulation eyewear, he could refine the font size and style selection within a small range, balancing readability and density of information. The design student observed that, although larger texts had initially been selected, the simulation effectively underscored the necessity of further font size adjustments to better accommodate presbyopic patients. Both designers expressed that the first-person-view simulation enriched their understanding and perception of how everyday objects might appear to their parents and design audiences. This enhanced understanding led to a refined bottle design featuring a larger size, pinchable bottom, and reduced neck and cap diameter for improved dosage control. The design student also highlighted a difference between monocular and binocular viewing, noting that vision seemed much clearer (“as usual”) when looking through the simulation eyewear with only one eye open. This difference could be leveraged to facilitate a smooth transition from normal vision (single-eye viewing) to a presbyopic simulation (dual-eye viewing) without removing the simulation eyewear during design concept evaluations.
4.7. Part III: Observations
Given the limited sample size ($N=2$), we frame these findings strictly as preliminary observations intended to guide future inquiry rather than as generalized evidence. However, the distinct design modifications produced by both the professional and the student (Figure 10) point toward three potential utilities of the OpticalAging approach:
@@ 4.7. Part III: Observations
    Facilitating Embodied Evaluation: The simulation appeared to bridge the gap between abstract guidelines and immediate perception. The specific adjustments to font size and bottle ergonomics were driven by the designers’ firsthand difficulty in reading and handling the prototypes. This aligns with Suri’s concept of ‘designing in a state of empathy’ (Suri, 2003), suggesting that tunable lenses could serve as a mechanism for generating embodied insight during the drafting phase.
@@
    Potential for Rapid Iteration: The participants’ ability to instantly toggle between simulated presbyopia and corrected vision seemed to support a tighter feedback loop. This observation suggests that such tools might enhance process efficiency, allowing designers to “spot check” accessibility constraints in real-time without the logistical overhead of recruiting testers for every minor iteration.
@@
    Perceptual Nuances in Usage: The feedback regarding monocular versus binocular viewing hints at opportunities for more sophisticated simulation interactions. Future work could explore how transitioning between these states might better mimic the fluctuating nature of visual fatigue.
Ultimately, this case study illustrates that while simulation cannot replicate the lived experience, context, or compensatory strategies of aging, it may function as a valuable design probe. It offers a “first line of defense” against inaccessible design, prompting immediate modifications (e.g., larger fonts, pinchable shapes) that might otherwise be overlooked until later testing stages. We emphasize that this tool is intended to complement, not replace, authentic user engagement with older adults.
5. Discussion
Our findings signal a paradigm shift in inclusive design methodology: moving from passive accessibility checklists to active embodied verification. Traditionally, accessibility guidelines function as external constraints—rules to be memorized and applied. However, as Hitchcock et al. argue, designers often disregard ergonomic data tables simply because they lack an intuitive grasp of the user’s actual constraints (Hitchcock et al., 2001). Our study supports this observation: designers using OpticalAging reported that the physical sensation of visual fatigue—specifically the soreness in their arms from holding objects at a distance—prompted immediate design changes (e.g., pinchable bottles) that abstract guidelines had failed to inspire.
This shift in perspective is further evidenced by how participants attributed their difficulties. Similar to Lavallière et al.’s findings with the AGNES suit, where participants attributed difficulty to the “store layout” rather than the simulation tool itself (Lavallière et al., 2017), our participants shifted their critique from their own vision to the design of the objects (e.g., “this font is too small” rather than “my eyes are blurry”). This indicates a successful shift toward inclusive design thinking, where the environment is viewed as the malleable variable rather than the user’s ability.
Finally, while we acknowledge the critique that increased understanding may stem from information exposure rather than simulation alone, the somatic data collected suggests otherwise. The physical fatigue reported by participants serves as a unique data point that verbal briefings cannot convey. By injecting the physiological constraints of aging directly into the designer’s visual stream, the system converts accessibility from a compliance task into an immediate, haptic reality.
Limitations
Several limitations should be considered when interpreting these findings: (1) Simulation Fidelity: As noted by critiques in disability studies (Bennett and Rosner, 2019; Tigwell, 2021), simulation risks oversimplification. Our system currently renders a uniform blur, which differs from the complex reality of aging vision that often includes ‘ghosting’, scattering, or astigmatism (Kleves, 2021; Fresina et al., 2012). Therefore, our tool represents an idealized model of presbyopia (pure accommodative loss) rather than a comprehensive replication of the condition.
(2) Study Scope and Generalizability: The user studies focused solely on the 50s simulation mode and a short duration. Exploring onset (40s) or advanced (60s+) modes over longitudinal periods is necessary to understand long-term adaptation. Furthermore, the exploratory case study ($N=2$) serves only as an illustrative workflow example; larger-scale validation with diverse design tasks is required to generalize the impact on professional practice.
@@ 6. Conclusion and Future Works
This paper presented OpticalAging, a real-time, optical see-through approach using tunable lenses to simulate the first-person visual perspective of presbyopia. Our findings, blending quantitative near-point validation with qualitative insights from designers, suggest that experiencing simulated presbyopic blur can significantly enhance understanding and awareness of perceptual challenges, potentially leading to more intuitive and user-centered designs for older adults. By providing a tool to help bridge the gap between abstract accessibility knowledge and embodied appreciation of perceptual challenges, this work offers a step towards more effective age-inclusive design practices.
@@
By shifting the focus from correction to simulation, we demonstrate how these powerful devices can serve not only patients but also the designers, educators, and engineers working to build a more accessible world. Ultimately, this research contributes more than a novel optical hardware setup; it proposes a new interface for intergenerational understanding. By shifting the focus of tunable lenses from vision correction to vision simulation, we demonstrate how wearable displays can serve not just patients, but the architects of our daily environments. OpticalAging transforms the designer’s own body into a high-fidelity proxy for the user, effectively reducing the cognitive distance between the ‘specification’ of a disability and the ‘sensation’ of living with it. While simulation can never fully replace the lived experience or compensatory strategies of older adults, providing designers with the ability to momentarily inhabit these perceptual constraints is a crucial step toward a world designed for the eyes of our future selves.
Acknowledgements.
This work was supported by JST Moonshot R&D Grant Number JPMJMS2012, JST ASPIRE Grant Number JPMJAP2327, and JST FOREST Grant Number JPMJFR206E; JSPS KAKENHI Grant Number JP20H05958 and JSPS Grant-in-Aid for Early-Career Scientists Grant Number JP25K21241; and JPNP23025 commissioned by the New Energy and Industrial Technology Development Organization (NEDO).
Appendix A Appendix